# DualReal: Adaptive Joint Training for Lossless Identity-Motion Fusion in Video Customization

### v1

@@ Abstract
Customized text-to-video generation with pre-trained large-scale models has recently garnered significant attention through focusing on identity and motion consistency. Existing works typically follow the isolated customized paradigm, where the subject identity or motion dynamics are customized exclusively. However, this paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrades. To address this, we introduce DualReal, a novel framework that, employs adaptive joint training to collaboratively construct interdependencies between dimensions. Specifically, DualReal is composed of two units: (1) Dual-aware Adaptation dynamically selects a training phase (i.e., identity or motion), learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid knowledge leakage; (2) StageBlender Controller leverages the denoising stages and Diffusion Transformer depths to guide different dimensions with adaptive granularity, avoiding conflicts at various stages and ultimately achieving lossless fusion of identity and motion patterns. We constructed a more comprehensive benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics. Project page: https://wenc-k.github.io/dualreal/
Figure 1: Generated customization results of our proposed novel paradigm DualReal. Given subject images and motion videos, DualReal generates high-quality customized identity and motion simultaneously, without compromising the consistency of either dimension.
@@ 1 Introduction
However, the existing isolated customized paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrades either motion coherence, subject fidelity, or both dimensions simultaneously. The reason is that the step-by-step diffusion video synthesis process itself, by nature, with different denoising steps dynamically reweights their spatiotemporal focus, i.e., progressively refines identity details across frames, enabling completely modeling through increasing denoising steps. The existing isolated customized paradigm violates this natural progression through indiscriminate dimensional over-specialize across all time steps, since they enforce uniform step sampling during motion/identity customized training, thereby inducing conflicting optimization trajectories between motion and identity accuracy. Consequently, existing methodologies inevitably cause mutual performance deterioration, that is, motion customization undermines the pre-trained video model’s inherent identity priors, or vice verse. As shown in Fig. 2, (1) adding the motion prior irreversibly reduced identity fidelity, indicating that isolated training fails to resolve dimensional conflicts during inference and leads to performance degradation; and (2) as motion training steps increased, optimal fidelity for different identities occurred unpredictably, suggesting that no universal number of training steps minimizes degradation. In summary, the existing paradigm does not meet the consistency and flexibility requirements of video customization tasks.
@@
Figure 2: Visual analysis of isolated training paradigm. We select different identities with the same motion pattern, fix the number of identity training steps, and gradually increase the number of motion training steps to achieve two-dimension customization. The red box marking the relative optimal position of the same identity’s fidelity. Experiments show that (1) adding motion prior significantly damages identity consistency; (2) We cannot find a universal step to minimize identity degradation from the positions of the red boxes for different identities.
@@
In this paper, we introduce DualReal, a novel framework that, for the first time, employs adaptive joint training to collaboratively construct interdependencies between identity and motion, which meets the consistency requirements of customized video generation in both identity and motion, as shown in Fig. 1. Technically, DualReal is composed of two complementary units: (1) Dual-aware Adaptation dynamically selects the appropriate training phase(i.e., either identity or motion) to learn the current information, guided by the other dimension prior. It also employs a regularization strategy to prevent dimensional knowledge leakage by blocking the parameter updates of the non-training dimension using gradient masking, thereby achieving effective joint training; (2) StageBlender Controller operates through coordinated utilization of denoising-stage progression and Diffusion Transformer (DiT) layer-depth variations during training. By adaptively allocating hierarchical focus (i.e., fine-grained adjustments for identity patterns and motion dynamics), it resolves dimensional competition across processing stages. This granularity-aware guidance ultimately achieves dimensional lossless fusion.
@@
Performance We constructed a more comprehensive evaluation benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our adaptive joint training framework.
@@ 2.1 Text-to-video Diffusion Models
Text-to-video generation aims to generate realistic videos based on prompts and has recently received growing attention [8, 1, 53, 30, 15, 44, 38, 11, 50, 52, 55, 45, 42, 28]. Current text-to-video generation architectures primarily fall into two categories[33, 29]. UNet-based video diffusion frameworks utilize hierarchical enc-dec with spatiotemporal learning [38, 11, 16, 3, 50, 44, 15], e.g., Video diffusion models [15] pioneered diffusion model applications in video generation through pixel-space video distribution modeling. Make-A-Video [38] and AnimateDiff [11] augment pretrained text-to-image models with motion modules. Recent advances in scalability drive the shift toward transformer-based architectures with joint spatiotemporal modeling [28, 52, 20, 26, 22], achieving revolutionary progress in video generation. Sora [28] introduces the diffusion-transformer framework, achieving cinematic-quality extended video synthesis with temporal stability. CogVideoX [52] introduces an expert transformer with adaptive LayerNorm for enhanced text-video feature fusion. While diffusion transformers exhibit strong generative capacities, their architectural constraints in spatiotemporal decoupling inherently limit dynamic concept embedding.
@@ 2.2 Generation Model Customization
In contrast to domain-agnostic generation frameworks, customized visual synthesis exhibits superior adaptability in addressing personalized visual requirements via parametric adaptation mechanisms[6, 18, 7, 12, 34, 35, 37, 39, 47, 9]. Textual inversion [9] aligns visual-textual semantics through text embedding optimization. Dreambooth [34] through full model fine-tuning of diffusion architectures to inject subject-specific priors. Building upon these foundational approaches, contemporary video customization research explores analogous methodologies[27, 5, 25, 13, 48, 49, 32, 2, 10]. MotionBooth [49] proposes comprehensive video diffusion model fine-tuning coupled with attention map manipulation for motion control during inference. DreamVideo [48] develops decoupled adapter training with joint inference mechanisms, coordinating subject customization and motion preservation during generation.
@@ 3 Methodology
Figure 3: Overall framework of DualReal. At each training step, we first dynamically select the training phase Z(i.e., identity or motion) to determine the data processing path. The specific data undergoes noise injection and combines with the text embeddings. StageBlender Controller governs two-dimension adapters’ contributions in Dual-Aware Block (DA-Block) through time-aware conditioning of current denoising step and fused feature representations. In DA-Block, the training-stage(i.e. Z) adapter learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid dimensional knowledge leakage, achieving joint training. Both branches engage in residual connections with DiT outputs.
Given a series of identity-specific images and motion sequences, DualReal through an innovative joint training framework, synthesizes coherent motion while preserving full-frame identity fidelity, as detailed in Fig. 3. During training, DualReal dynamically switches optimization focus between identity and motion, adjusting the module parameters through specific guidance. To address the key challenges: (1) enabling joint identity-motion interaction modeling in unified parameter spaces. (2) mitigating attribute leakage risks in alternating training modes. We propose the Dual-aware Adaptation architecture with a complementary regularization strategy in Sec. 3.2, forming an integrated framework for parameter-shared adaptation. Moreover, we propose StageBlender Controller, which leverages the denoising stages and DiT depths to guide different modes with adaptive granularity, avoiding dimensional conflicts at various stages and ultimately achieving lossless fusion of identity and motion in Fig. 5.
@@ 3.1 Preliminary
DiT-based Models. Most DiT-based models [52, 28, 22, 26] process concatenated conditioning prompt and spatiotemporal visual tokens through transformer layers, establishing multi-modal coupling between text-guided semantic contexts and visual representations in latent space. Despite achieving a remarkable capabilities in generic generation scenarios, this architectural fundamentally constrains conventional personalization frameworks that demand decoupled control along spatial-temporal axes [46, 49].
@@ 3.2 Dual-aware Adaptation
To achieve joint training of identity and motion and resolve dimensional conflicts, we innovatively proposed Dual-aware Adaptation, which leverages the prior from one dimension to guide the training of the other while preventing information leakage through a regularization strategy, as shown in the lower half of Fig. 3.
Joint Identity-Motion Optimization. Different from some approaches that fine-tune the whole diffusion model  [34], DualReal first dynamically select the training dimension (i.e., motion-focused or identity-focused) with the predefined hyperparameter ratio before each denoising iteration, the corresponding data is then sent to the DiT. The input of $i$-th block is the joint feature $f_{\text{in}}^{i}=[f_{\text{text}}^{i},f_{\text{viusal}}^{i}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, where $n_{t}$ , $n_{v}$ represent the number of text and visual tokens respectively, adapters employ the bottleneck architecture with skip connections  [48], which can be expressed as:
@@
where the activation function $\sigma$ corresponds to GELU  [14], $\mathbf{W}$ and $\mathbf{W^{{}^{\prime}}}$denote the identity-preserving and motion-specific linear projection weights respectively, both operating on hidden dimension $d$. The weight $\mathbf{W}_{\text{{cond}}}\in\mathbb{R}^{e\times c}$ of conditional linear maps reference image embedding to the latent space [48].
Through StageBlender Controller constraint (Fig. 5) , the motion adapter outputs are scaled by weight coefficient $\omega_{i}$ , with identity outputs weighted by the complementary coefficient ($1-\omega_{i}$). The modulated features are aggregated into the output of DiT blocks through residual connections, above process can be formulated as:
@@
where $f_{\text{dit}}^{i}$ denoting the output of the $i$-th DiT layer and $\hat{f}_{\text{out}}^{i}$ indicating the aggregated output of final block. This parametric constraint intrinsically balances feature contributions across blocks and denoising stages through complementary scaling, while structurally enforcing dedicated attention to either identity preservation or motion dynamics during adaptation.
Regularization Strategy. A critical challenge in joint dimensional training arises from the significant distribution shift across different training dimension, where unconstrained optimization usually causes destructive interference between cross-dimension knowledge observed in previous work [51, 49]. For example, fine-tuning the motion adapter with static images during the temporal training phase irreversibly degrades its dynamic generation capability, with analogous effects occurring during identity adaptation. To resolve this, we employ the Regularization with the gradient mask M, to activate only the corresponding adapter parameters based on the selected stage Z, optimizing motion coherence(i.e., through motion adapter), or preserving identity consistency(i.e., through identity adapter), which can be formulated as:
$\theta^{(t+1)}=\theta^{(t)}-\bf M\odot\nabla_{\theta}\mathcal{L}\ ,$ (4)
${\bf M}={\bf Z}\cdot M_{m}+(1-{\bf Z})\cdot M_{i}\ .$ (5)
@@
where $\mathcal{L}$ denotes the video diffusion reconstruction loss, Z is the binary selector, and $\theta_{m}/\theta_{i}$ represent the motion/identity adapter parameters respectively.
Simultaneously, we keep frozen adapter in a waiting state to inform modal expertise for active adapter during forward propagation, enabling two dimensional features referencing within current data streams. The features from frozen adapter act as intrinsic regularization to constrain dimension overfitting, thereby facilitating mutual reference learning without interference.
@@ 3.3 StageBlender Controller
Figure 4: Illustration of proposed StageBlender Controller, which employs Adaptive LayerNorm mechanism modulates text-visual feature based on timestep-conditional embeddings, then maps the feature to multiple groups after residual gated connections. These scaled weight are subsequently routed to their respective DA-Blocks for processing.
@@
Furthermore, in order to resolves dimensional competition across processing stage, we propose the StageBlender Controller that governs dimensional contributions through time-aware conditioning of block-level scaling coefficients, which empowers the DA-Block to adaptively allocate specific dimension shares (i.e., achieve granularity decoupling) through the mechanism detailed in Fig. 4. Specifically, this module dynamically generates multiple sets of scaling weights according to denoising timestep embedding and the fused text-visual features. For the input feature $f_{\text{in}}=[f_{\text{text}}^{1},f_{\text{viusal}}^{1}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, the processing flow first extracts salient features through pooling, then adaptively modulates them via DiT Adaptive LayerNorm  [52] with injected timestep embeddings $\bf t$. This operation can be formulated as:
@@
where the $\mathbf{W}\in\mathbb{R}^{c\times\mathbf{t}_{\text{dim}}}$ denotes the weight matrix with $\mathbf{t}_{\text{dim}}$ as the channel dimension of timestep embedding, and $\alpha$, $\beta$ are defined as:
@@
The computed weights coefficients are then integrated to enable gated fusion between timestep and visual text tokens, as formulated below:
@@
where $\mathbf{L}$ corresponds to the depth of the transformer block, $\mathbf{n}$ specifies the cardinality of the disentangled weight groups and $\mathbf{\Gamma}$ is the projection operator: $\mathbb{R}^{\mathbf{L}}\to\mathbb{R}^{\mathbf{n}}$. Each group sequentially governs designated DA-Block layers through parameter assignment.

### v2

@@ Abstract
Customized text-to-video generation with pre-trained large-scale models has recently garnered significant attention by focusing on identity and motion consistency. Existing works typically follow the isolated customized paradigm, where the subject identity or motion dynamics are customized exclusively. However, this paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrade. To address this, we introduce DualReal, a novel framework that employs adaptive joint training to construct interdependencies between dimensions collaboratively. Specifically, DualReal is composed of two units: (1) Dual-aware Adaptation dynamically switches the training step (i.e., identity or motion), learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid knowledge leakage; (2) StageBlender Controller leverages the denoising stages and Diffusion Transformer depths to guide different dimensions with adaptive granularity, avoiding conflicts at various stages and ultimately achieving lossless fusion of identity and motion patterns. We constructed a more comprehensive evaluation benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion metrics. Page: https://wenc-k.github.io/dualreal-customization
Figure 1: Generated customization results of our proposed novel paradigm DualReal. Given identity images and motion videos, DualReal generates high-quality customized identity and motion simultaneously, without compromising the consistency of either dimension.
@@ 1 Introduction
However, the existing isolated customized paradigm completely ignores the intrinsic mutual constraints and synergistic interdependencies between identity and motion, resulting in identity-motion conflicts throughout the generation process that systematically degrade either motion coherence, subject fidelity, or both dimensions simultaneously. The reason is that the step-by-step diffusion video synthesis process itself, by nature, with different denoising steps dynamically reweights their spatiotemporal focus, i.e., progressively refines identity details across frames, enabling complete modeling through increasing denoising steps. The existing isolated customized paradigm violates this natural progression through indiscriminate dimensional over-specialization across all time steps, since they enforce uniform step sampling during motion/identity customized training, thereby inducing conflicting optimization trajectories between motion and identity accuracy. Consequently, existing methodologies inevitably cause mutual performance deterioration, that is, motion customization undermines the pre-trained video model’s inherent identity priors, or vice versa. As shown in Fig. 2, (1) adding the motion prior irreversibly reduced identity fidelity, indicating that isolated training fails to resolve dimensional conflicts during inference and leads to performance degradation; and (2) as motion training steps increased, optimal fidelity for different identities occurred unpredictably, suggesting that no universal number of training steps minimizes degradation. In summary, the existing paradigm does not meet the consistency and flexibility requirements of video customization tasks.
@@
Figure 2: Visual analysis of isolated training paradigm. We select different identities with the same motion pattern, fix the number of identity training steps, and gradually increase the number of motion training steps to achieve two-dimensional customization. The red box marking the relative optimal position of the same identity’s fidelity. Experiments show that (1) adding motion prior significantly damages identity consistency; (2) We cannot find a universal step to minimize identity degradation from the positions of the red boxes for different identities.
@@
In this paper, we introduce DualReal, a novel framework that, for the first time, employs adaptive joint training to collaboratively construct interdependencies between identity and motion, which meets the consistency requirements of customized video generation in both identity and motion, as shown in Fig. 1. Technically, DualReal is composed of two complementary units: (1) Dual-aware Adaptation dynamically switches the training step (i.e., either identity or motion) to learn the current information, guided by the other dimension prior. It also employs a regularization strategy to prevent dimensional knowledge leakage by blocking the parameter updates of the non-training dimension using gradient masking, thereby achieving effective joint training; (2) StageBlender Controller operates through coordinated utilization of denoising-stage progression and Diffusion Transformer (DiT) layer-depth variations during training. By adaptively allocating hierarchical focus (i.e., fine-grained adjustments for identity patterns and motion dynamics), it resolves dimensional competition across processing stages. This granularity-aware guidance ultimately achieves dimensional lossless fusion.
@@
Performance We constructed a more comprehensive evaluation benchmark than existing methods. The experimental results show that DualReal improves CLIP-I and DINO-I metrics by 21.7% and 31.8% on average, and achieves top performance on nearly all motion quality metrics, demonstrating the efficiency of our framework.
@@ 2.1 Text-to-video Diffusion Models
Recent advances in generative models have significantly improved the quality and versatility of synthetic content[20, 19, 16, 42, 59, 9]. DQVAE [20] generates images autoregressively in a more effective coarse-to-fine order. Text-to-video generation aims to generate realistic videos based on prompts and has recently received growing attention [1, 57, 34, 48, 12, 54, 56, 49, 46, 32]. Current text-to-video generation architectures primarily fall into two categories [37, 33]. UNet-based video diffusion frameworks utilize hierarchical enc-dec with spatiotemporal learning [42, 12, 17, 3, 54, 48, 16], e.g., Video diffusion models [16] pioneered diffusion model applications in video generation through pixel-space video distribution modeling. Make-A-Video [42] and AnimateDiff [12] augment pretrained text-to-image models with motion modules. Recent advances in scalability drive the shift toward transformer-based architectures with joint spatiotemporal modeling [32, 56, 23, 30, 25], achieving revolutionary progress in video generation. Sora [32] introduces the diffusion-transformer framework, achieving cinematic-quality extended video synthesis with temporal stability. CogVideoX [56] introduces an expert transformer for enhanced text-video feature fusion. While diffusion transformers exhibit strong generative capacities, their architectural constraints in spatiotemporal decoupling inherently limit dynamic concept embedding.
@@ 2.2 Generation Model Customization
Generation Model Customization has emerged as a pivotal strategy[41, 43, 7, 21, 29]. In contrast to domain-agnostic generation frameworks, customized visual synthesis exhibits superior adaptability in addressing personalized visual requirements via parametric adaptation mechanisms[6, 8, 13, 38, 39, 51, 10, 20]. Textual inversion [10] aligns visual-textual semantics through text embedding optimization. Dreambooth [38] through full model fine-tuning of diffusion architectures to inject subject-specific priors. RealCustom[21, 29] disentangles similarity from controllability by precisely limiting subject influence to relevant parts only. Building upon these foundational approaches, contemporary video customization research explores analogous methodologies[31, 5, 28, 14, 52, 53, 36, 2, 11]. MotionBooth [53] proposes a comprehensive video diffusion model fine-tuning coupled with attention map manipulation for motion control during inference. DreamVideo [52] develops decoupled adapter training with joint inference mechanisms, coordinating subject customization and motion preservation during generation.
@@ 3 Methodology
Figure 3: Overall framework of DualReal. At each training step, we first dynamically switch the training step $Z$ (i.e., identity or motion) to determine the data processing path. The specific data undergoes noise injection and combines with the text embeddings. StageBlender Controller governs two-dimensional adapters’ contributions in Dual-Aware Block (DA-Block) through time-aware conditioning of current denoising step and fused feature representations. In DA-Block, the training-stage adapter learns the current information guided by the frozen dimension prior, and employs a regularization strategy to avoid dimensional knowledge leakage, achieving joint training. Both branches engage in residual connections with DiT outputs.
Given a series of identity-specific images and motion sequences, DualReal through an innovative joint training framework, synthesizes coherent motion while preserving full-frame identity fidelity, as detailed in Fig. 3. During training, DualReal dynamically switches optimization focus between identity and motion, adjusting the module parameters through specific guidance. To address the key challenges: (1) enabling joint identity-motion interaction modeling in unified parameter spaces. (2) mitigating attribute leakage risks in alternating training dimensions. We propose the Dual-aware Adaptation architecture with a complementary regularization strategy in Sec. 3.2, forming an integrated framework for parameter-shared adaptation. Moreover, we propose StageBlender Controller, which leverages the denoising stages and DiT depths to guide different dimensions with adaptive granularity, avoiding conflicts at various stages and ultimately achieving high-fidelity fusion of identity and motion in Fig. 5.
@@ 3.1 Preliminary
DiT-based Models. Most DiT-based models [56, 32, 25, 30] process concatenated conditioning prompt and spatiotemporal visual tokens through transformer layers, establishing multi-modal coupling between text-guided semantic contexts and visual representations in latent space. Despite achieving remarkable capabilities in generic generation scenarios, this architecture fundamentally constrains conventional personalization frameworks that demand decoupled control along spatial-temporal axes [50, 53].
@@ 3.2 Dual-aware Adaptation
To achieve joint training of identity and motion while resolving dimensional conflicts, we innovatively proposed Dual-aware Adaptation, which leverages the prior from one dimension to guide the training of the other while preventing information leakage through a regularization strategy, as shown in the lower half of Fig. 3.
Joint Identity-Motion Optimization. Different from some approaches that fine-tune the whole diffusion model  [38], DualReal first dynamically switches the training step (i.e., motion-focused or identity-focused) with the predefined hyperparameter ratio before each denoising iteration; the corresponding data is then sent to the DiT. The input of $i$-th block is the joint feature $f_{\text{in}}^{i}=[f_{\text{text}}^{i},f_{\text{visual}}^{i}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, where $n_{t}$, $n_{v}$ represent the number of text and visual tokens respectively. The adapters employ the bottleneck architecture with skip connections  [52]:
@@
where the activation function $\sigma$ corresponds to GELU  [15], $\mathbf{W}$ and $\mathbf{W^{{}^{\prime}}}$denote the identity and motion linear projection weights, respectively, both operating on hidden dimension $d$. The weight $\mathbf{W}_{\text{{cond}}}\in\mathbb{R}^{e\times c}$ of conditional linear maps reference image embedding to the latent space [52].
Through StageBlender Controller constraint (Fig. 5), the motion adapter outputs are scaled by weight coefficient $\omega_{i}$, with identity outputs weighted by the complementary coefficient ($1-\omega_{i}$). The modulated features are aggregated into the output of DiT blocks through residual connections. The above process can be formulated as:
@@
where $f_{\text{dit}}^{i}$ denotes the output of the $i$-th DiT layer and $\hat{f}_{\text{out}}^{i}$ indicates the aggregated output of the final block. This parametric constraint intrinsically balances feature contributions across blocks and denoising stages, while structurally enforcing dedicated attention to either identity preservation or motion dynamics during adaptation.
Regularization Strategy. A critical challenge in joint dimensional training arises from the significant distribution shift across different training dimensions, where unconstrained optimization usually causes destructive interference between cross-dimension knowledge observed in previous work [55, 53]. For example, fine-tuning the motion adapter with static images during the temporal training phase irreversibly degrades its dynamic generation capability, with analogous effects occurring during identity adaptation. To resolve this, we employ regularization with the gradient mask $M$ to activate only the corresponding adapter parameters based on a binary selector variable $Z\in\{0,1\}$, optimizing motion coherence(i.e., through motion adapter), or preserving identity consistency(i.e., through identity adapter), which can be formulated as:
$\theta^{(t+1)}=\theta^{(t)}-M\odot\nabla_{\theta}\mathcal{L}\ ,$ (4)
$M=Z\cdot M_{m}+(1-Z)\cdot M_{i}\ .$ (5)
@@
where $\mathcal{L}$ denotes the video diffusion reconstruction loss. The adapter parameters $\theta$ are split into motion ($\theta_{m}$) and identity ($\theta_{i}$) components using binary masks $M_{m}$ and $M_{i}$.
Simultaneously, we keep a frozen adapter in a waiting state to inform modal expertise for active adapters during forward propagation, enabling two-dimensional features referencing within current data streams. The features from the frozen adapter act as intrinsic regularization to constrain dimension overfitting, thereby facilitating mutual reference learning without interference.
@@ 3.3 StageBlender Controller
Figure 4: Illustration of proposed StageBlender Controller, which employs an Adaptive LayerNorm mechanism that modulates text-visual feature based on timestep-conditional embeddings, then maps the feature to multiple groups after residual gated connections. These scaled weights are subsequently routed to their respective DA-Blocks for processing.
@@
Furthermore, in order to resolve dimensional competition across the processing stage, we propose the StageBlender Controller that governs dimensional contributions through time-aware conditioning of block-level scaling coefficients, which empowers the DA-Block to adaptively allocate specific dimension shares (i.e., achieve granularity decoupling) through the mechanism detailed in Fig. 4. Specifically, this module dynamically generates multiple sets of scaling weights according to denoising timestep embedding and the fused text-visual features. For the input feature $f_{\text{in}}=[f_{\text{text}}^{1},f_{\text{visual}}^{1}]\in\mathbb{R}^{B\times(n_{t}+n_{v})\times c}$, the processing flow first extracts salient features through pooling, then adaptively modulates them via DiT Adaptive LayerNorm  [56] with injected timestep embeddings $t$. This operation can be formulated as:
@@
where the $\mathbf{W}\in\mathbb{R}^{c\times t_{\text{dim}}}$ denotes the weight matrix with $t_{\text{dim}}$ as the channel dimension of timestep embedding, and $\alpha$, $\beta$ are defined as:
@@
The computed weight coefficients are then integrated to enable gated fusion between the timestep and visual text tokens, as formulated below:
@@
where $\mathbf{\Gamma}$ is the projection operator: $\mathbb{R}^{L}\to\mathbb{R}^{n}$. The $L$ denotes the DiT block depth, and $n$ specifies the number of disentangled weight groups; each group then sequentially controls its assigned layers via parameter assignment.